# AdaptGCD: Multi-Expert Adapter Tuning for Generalized Category Discovery

### v1

@@ 4.2 Comparison with State-of-the-Arts
Evaluation on generic image recognition datasets. AdaptGCD is compared with the previous advanced methods and some concurrent works (i.e., DCCL DBLP:conf/cvpr/PuZS23 , PromptCAL DBLP:conf/cvpr/promptcalgcdZhangKSNCK23 , SimGCD DBLP:conf/iccv/simgcdWenZQ23 and SPTNet wang2024sptnet ). The results are shown in Table 1. It is observed that our AdaptGCD manifests superior performance across three general benchmarks. Compared to the baseline SimGCD, AdaptGCD achieves improvements of 0.8%, 3.9%, 3.1% for “All” classes on three datasets. When juxtaposed with other concurrent methods, our method advances on all the metrics except for the “Old” classes on CIFAR10. On CIFAR10 datasets, AdaptGCD prioritizes a balance between the old and new categories so that it outperforms the GCD by 10.8% in “New” classes, at a cost of 2% dip in the “Old” classes. It is also noted that the most advanced prompt-based method, SPTNet, underperforms its baseline SimGCD on several metrics on CIFAR10 and CIFAR100 datasets since the low-resolution images hinder the utility of spatial prompts. The promising performance achieved by our AdaptGCD further validates that the adapter-like methods could exhibit excellent adaptability to data with varying qualities than the prompt-like rivals.
@@
Evaluation on fine-grained datasets. Table 2 presents the results on fine-grained datasets including the semantic shift benchmark and the long-tail dataset Herbarium19. Due to the large intra-class and small inter-class variations for data, it is more difficult to discover fine-grained categories than the generic classes. Despite the difficulty, AdaptGCD shows consistently superior performance on the CUB-200, Scars and Herbarium19 datasets, achieving an average improvement of 3% on “All” classes in comparison with the state-of-the-art results. Even within the Aircraft dataset, it achieves the second best scores with only 1.4% lag on “All” classes. These observations indicate the efficacy of the proposed method across fine-grained datasets.
4.3 Ablation Study
@@
To investigate the impact of different modules, we conduct the the extensive ablation studies on the semantic shift benchmark following wang2024sptnet ; DBLP:conf/cvpr/promptcalgcdZhangKSNCK23 . The results are reported in Table 3.
Table 3: Ablation study on the semantic shift benchmark
@@
All Old New All Old New All Old New
(1) SimGCD - - $\mathcal{L}_{gcd}$ 60.3 65.6 57.7 53.8 71.9 45.0 54.2 59.1 51.8
(2) SimGCD + MEA 1 64 $\mathcal{L}_{gcd}$ 67.5 75.3 63.5 59.8 77.4 51.2 55.2 65.0 50.3
(3) SimGCD + MEA 1 512 $\mathcal{L}_{gcd}$ 68.0 76.1 63.9 61.3 77.5 53.5 57.7 66.9 53.1
(4) SimGCD + MEA 8 64 $\mathcal{L}_{gcd}$ 68.5 74.9 65.3 61.4 79.4 52.7 57.3 65.3 53.2
(5) SimGCD + MEA 8 64 $\mathcal{L}_{gcd}+\mathcal{L}_{bl}$ 68.6 75.0 65.4 62.0 80.9 52.8 57.7 65.3 53.9
(6) SimGCD + MEA 8 64 $\mathcal{L}_{gcd}+\mathcal{L}_{bl}+\mathcal{L}_{pbl}$ 68.8 74.5 65.9 62.7 80.6 54.0 57.9 65.2 54.3
Influence of adding adapter structure. Firstly, we have validated the role of a single adapter as illustrated in (2). Given that the bottleneck dimension is set as 64, the count of learnable parameters in the backbone is less than 0.8M, which is significantly less than the number of parameters in the last block (approximately 7M). Despite the less trainable parameters in the backbone, involving an adapter still achieves remarkable accuracy boost of 7.2%, 6.0% and 1.0% for “All" classes on the CUB-200, Scars and Aircraft datasets. These findings affirm the adapter’s efficacy for GCD task.
Influence of multi-expert adapter and route assignment constraint. To validate the effectiveness of the multi-expert adapter (MEA), the method implemented with MEA (in (4)) is compared with two alternate single adapter approaches. One incorporates an equivalent bottleneck dimension, as shown in (2), while the other uses the setup with a comparable parameter count, denoted in (3). With an identical bottleneck dimension, the results reveal that the MEA outperforms the single expert method, thereby establishing the MEA’s superiority. On the other hand, when the number of learnable parameters is similar (see (3) vs. (4)), the MEA does not show obvious advantages. Nevertheless, with the assistance of route assignment constraints (see (5),(6)), the methods incorporating MEA surpass the single adapter method (3) in terms of “All" classes across three datasets. This statement emphasizes the significant role that the route assignment constraint holds in the MEA training process. It is worth noting that the improvements derived from the application of MEA are not only due to the increased number of learnable parameters but also contribute to the separation of different data via route allocations, enabled by the constraint. It’s also observed that the $\mathcal{L}_{pbl}$ enhances the performance of “New” classes at the expense of a slight reduction in “Old” classes (see (6) vs. (5)). This observation also validates our motivation to alleviate the bias towards old classes.
4.4 Discussion
Figure 2: Sensitivity analysis on three critical hyper-parameters. (a) bottleneck dimension $\hat{d}$, (b) the number of adapted blocks $P$, (c) expert count $T$.
Sensitivity of three key hyper-parameters of the adapter. We evaluated the impact of three key hyper-parameters of the adapter: bottleneck dimension $\hat{d}$, number of adapted blocks $P$ and expert count $T$, and show the results in Figure 2. First, we conduct experiments with models employing single adapters across different $\hat{d}$ and report the results in Figure 2 (a). Remarkably, when the bottleneck dimension is relatively small, there is a noticeable rise in performance with the increase in the bottleneck dimension. As the dimension further elevates, performance fluctuates within a certain range due to the overfitting caused by the excessive amount of parameters. A similar trend is also observed as the increase in the number of adapted blocks $P$ in Figrue 2 (b), i.e., as $P$ is large enough, inserting more adapters in the bottle blocks yields additional resources but limited performance gain. It also validates the effectiveness of only inserting adapters in parts of blocks. For the high performance of a single adapter, the impact of $T$ in Figure 2 (c) is not so noticeable, but the increment in expert count still results in a performance boost for “All” classes, further corroborating the effectiveness of multi-expert adapters.
Figure 3: Attention visualization for 12 heads in the last blocks of backbone on CUB-200. “DINO” represents the results from pretrained backbone without additional training while “SimGCD” and “AdaptGCD” are the those from models trained via SimGCD and AdaptGCD.
Visualization of the attended maps. To better appreciate the advantages of the AdaptGCD, we extracted the attention maps corresponding to the [CLS] token from different heads in the last block of the ViT, highlighting the top 10 most notable patches in red for each head in Figure 3. Upon observation, it is evident that compared to SimGCD and DINO, AdaptGCD focuses more on the foreground and exhibits more varied attention regions in different heads (in heads 3/7/11/12), showing the superior adaptability of AdaptGCD to the CUB-200 datasets. It is also worth noting that in regions like heads 3/7/11, SimGCD fails to inherit the correct attention on the foreground from the DINO model while AdaptGCD refines the attention regions based on the DINO model. These findings further demonstrate the motivation to protect the pretrained information for generalization.
Figure 4: (a) Visualization of the route assignment probability for samples predicted as old and new classes in 8 experts. (b) Confusion maps for the old and new classes.
Visualization of route assignment probability. To verify the role of the route assignment constraint, we also visualize the assignment probabilities for samples predicted as old and new classes under the original SimGCD loss, the loss with $\mathcal{L}_{bl}$, and the loss with $\mathcal{L}_{bl}+\mathcal{L}_{pbl}$ on Aircraft dataset in Figure 4 (a). Without any additional loss, the “winner-takes-all” phenomenon occurs and the imbalance is greatly alleviated via $\mathcal{L}_{bl}$. Note that a significant portion of samples predicted as old or new classes are distributed to their corresponding experts with the addition of $\mathcal{L}_{bl}$ and $\mathcal{L}_{pbl}$. It validates the effectiveness of $\mathcal{L}_{pbl}$ to separate the data from old and new classes on the route level.
@@ 4.4 Discussion
Confusion matrix of the model trained with balanced loss. In addition, to further investigate the role of the route assignment constraint in mitigating the bias towards old classes, we also visualized the confusion matrix of the corresponding method in Figure 4 (b). It is observed that with the constraint, the amount of new-class data misclassified as the old classes decreases, thus substantiating the effectiveness of the constraint in alleviating the preference for the old classes.
5 Conclusion
@@
In this study, we propose a novel AdaptGCD in generalized category discovery, which is the pioneering work to merge adapter tuning into the GCD framework. It integrates learnable adapters with the fixed backbone to adapt to the downstream GCD task as well as preserve the pretrained priority. Additionally, considering the imbalance between the supervision of the old and new classes, we present a multi-expert adapter structure and a route assignment constraint to separate the data from old and new classes into different expert groups, further reducing mutual interference. The efficacy of the proposed strategy is demonstrated in 7 datasets, including 3 generic datasets 3 fine-grained datasets, and a long-tail dataset. Looking forward to future work, a promising direction is to explore more effective fine-tuning strategies in the GCD tasks, including the LoRA DBLP:conf/iclr/HuSWALWWC22 , RepAdapter DBLP:journals/corr/abs-2302-08106 , or the combination of different strategies like GLoRA DBLP:journals/corr/gloraabs-2306-07967 . In terms of limitations, the proposed method introduces additional parameters, which consume extra computational cost during training and inference.

### v2

@@ V-B Comparison with State-of-the-Arts
1) Evaluation on Generic Image Recognition Datasets: AdaptGCD is compared with the previous advanced methods, including the methods based on partial tuning (ORCA [12], GCD [13], SimGCD [55], DCCL [53], CMS [78], AMEND [79], etc.) and those based on prompt tuning (PromptCAL [21], GCA+PromptCAL [80] and SPTNet [14]). The results on the generic image recognition benchmark are shown in Table III. It is observed that our AdaptGCD manifests superior performance across three general benchmarks. Compared to the baseline SimGCD, the AdaptGCD, which is equipped with fewer tunable parameters, achieves improvements of 0.8%, 3.9%, and 3.4% for “All” classes on three datasets, demonstrating the effectiveness of adapter tuning in striking the balance between the generalization and adaptability. When juxtaposed with other concurrent methods, our method advances on almost all the metrics. On CIFAR10 datasets, AdaptGCD prioritizes a balance between the old and new categories so that it outperforms the GCD by 10.8% in “New” classes, at a cost of 2.4% dip in the “Old” classes. Note that the most advanced prompt-based method SPTNet underperforms its baseline SimGCD on several metrics on CIFAR10 and CIFAR100 datasets since the low-resolution images hinder the utility of spatial prompts. However, poor image quality exerts a minimal impact on the proposed method. The promising performance achieved by our AdaptGCD validates that the adapter-like methods could exhibit excellent adaptability to data with varying image qualities than the prompt-like rivals. Furthermore, as a plug-in module, it also improves the performance of the existing method SPTNet.
@@
2) Evaluation on Fine-Grained Datasets: Table IV presents the results on fine-grained datasets including the semantic shift benchmark and the long-tailed dataset Herbarium19. Due to the large intra-class and small inter-class variations for data, it is more difficult to discover fine-grained categories than the generic classes. Despite the difficulty, AdaptGCD shows consistently superior performance on the CUB-200, Scars and Herbarium19 datasets, achieving an average improvement of 3% on “All” classes in comparison with the advanced results from SPTNet due to the excellent adaptability of adapter tuning. Even within the Aircraft dataset, it achieves the second best scores with only 1.4% lag on “All” classes. These observations indicate the efficacy of the proposed method across fine-grained datasets. Additionally, after incorporating the SPTNet, the proposal further achieves better performance on the ’All’ classes across the three datasets. This highlights the scalability of the proposed module and its potential to function as a plug-and-play component.
TABLE V: Ablation study on the semantic shift benchmark and generic image recognition benchmark. $T$ is the number of experts while $\hat{d}$ indicates the dimension number of each expert.
Index Method $T$ $\hat{d}$ Loss Semantic Shift Benchmark Generic Image Recognition Benchmark
CUB-200 Scars Aircraft CIFAR100 CIFAR10 Imagenet100
All Old New All Old New All Old New All Old New All Old New All Old New
(1) SimGCD - - $\mathcal{L}_{sgcd}$ 60.3 65.6 57.7 53.8 71.9 45.0 54.2 59.1 51.8 80.1 81.2 77.8 97.1 95.1 98.1 83.0 93.1 77.9
(2) AdaptGCD 1 64 $\mathcal{L}_{sgcd}$ 67.5 75.3 63.5 59.8 77.4 51.2 55.2 65.0 50.3 80.7 85.6 70.7 97.9 96.0 98.9 84.6 95.1 79.3
(3) AdaptGCD 1 512 $\mathcal{L}_{sgcd}$ 68.0 76.1 63.9 61.3 77.5 53.5 57.7 66.9 53.1 81.9 87.9 69.7 97.9 95.8 99.0 85.0 95.1 80.0
(4) AdaptGCD 8 64 $\mathcal{L}_{sgcd}$ 68.5 74.9 65.3 61.4 79.4 52.7 57.3 65.3 53.2 82.6 87.1 73.6 97.8 95.2 99.1 85.4 95.1 80.6
(5) AdaptGCD 8 64 $\mathcal{L}_{sgcd}+\mathcal{L}_{ba}$ 68.6 75.0 65.4 62.0 80.9 52.8 57.7 65.3 53.9 83.5 87.2 76.0 97.8 95.3 99.1 86.0 95.1 81.4
(6) AdaptGCD 8 64 $\mathcal{L}_{sgcd}+\mathcal{L}_{ba}+\mathcal{L}_{cba}$ 68.8 74.5 65.9 62.7 80.6 54.0 57.9 65.2 54.3 84.0 86.2 79.7 97.9 95.5 99.0 86.4 94.9 82.1
V-C Ablation Study
@@
To investigate the impact of different modules, we conduct extensive ablation studies on the semantic shift benchmark and generic image recognition benchmark following [14, 21]. The results are reported in Table V.
1) Influence of Adding Adapter Structure: Firstly, we have validated the role of a single adapter as illustrated in (2). Given that the bottleneck dimension is set as 64, the count of learnable parameters in the backbone is less than 0.8M, which is significantly less than the number of parameters in the last block (approximately 7M). Despite the less trainable parameters in the backbone, the adapter-based method still achieves a remarkable accuracy boost of 7.2%, 6.0%, 1.0%, 0.6%, 0.8% and 1.6% for “All” classes on the CUB-200, Scars, Aircraft, CIFAR100, CIFAR10 and Imagenet100 datasets. These findings affirm the adapter’s efficacy in GCD tasks in preserving the general knowledge and enhancing the adaptability to GCD task.
2) Influence of Multi-Expert Adapter and Route Assignment Constraint: To validate the effectiveness of the multi-expert adapter (MEA), the method implemented with MEA (in (4)) is compared with two alternate single adapter approaches. One incorporates an equivalent bottleneck dimension, as shown in (2), while the other uses the setup with a comparable parameter count, denoted in (3). With an identical bottleneck dimension, the results reveal that the MEA outperforms the single expert method, thereby establishing the MEA’s superiority. On the other hand, when the number of learnable parameters is similar (see (3) vs. (4)), the MEA does not show obvious advantages. Nevertheless, with the assistance of route assignment constraints (see (5),(6)), the methods incorporating MEA surpass the single adapter method (3), in terms of “All” classes across six datasets. This statement emphasizes the significant role that the route assignment constraint holds in the MEA training process. Actually, on average across three fine-grained and three general datasets, the two assignment losses $\mathcal{L}_{cba}$ and $\mathcal{L}_{ba}$ improve all class accuracy by 0.7% and 0.8%. It is further demonstrated that the improvements derived from the application of MEA are not only due to the increased number of learnable parameters but also contribute to the separation of different data via route allocations, enabled by the constraint. Moreover, the two losses enhance the performance of “New” classes by an average of 1% and 2.5% across fine-grained and general datasets, albeit at the expense of a slight reduction in “Old” classes (see (6) vs. (4)). This observation further validates our motivation to alleviate the bias towards old classes. More analysis about MEA is presented in the discussion section.
Figure 3: Sensitivity analysis on three critical hyperparameters. (a) bottleneck dimension $\hat{d}$, (b) the number of adapted blocks $P$, (c) expert count $T$.
V-D Sensitivity Analysis
1) Bottleneck Dimension, Number of Adapter Blocks and Expert Count: We evaluated the impact of three key hyper-parameters of the adapter: bottleneck dimension $\hat{d}$, number of adapted blocks $P$ and expert count $T$, and show the results in Fig. 3. First, we conduct experiments with models employing single adapters across different $\hat{d}$ and report the results in Fig. 3 (a). Remarkably, when the bottleneck dimension is relatively small, there is a noticeable rise in performance with the increase in the bottleneck dimension. As the dimension further elevates, performance fluctuates within a certain range due to overfitting caused by the excessive amount of parameters. A similar trend is also observed as the increase in the number of adapted blocks $P$ in Fig. 3 (b), i.e., as $P$ is large enough, inserting more adapters in the blocks yields additional resources but limited performance gain. It also validates the effectiveness of only inserting adapters in parts of blocks. For the high performance of a single adapter, the impact of $T$ in Fig. 3 (c) is not so noticeable, but the increment in expert count still results in a performance boost for “All” classes, further corroborating the effectiveness of multi-expert adapters.
2) Expert Assignment: To verify the impact of expert distribution, we conducted experiments with different numbers of new and old experts on the Scars and Aircraft datasets, with results shown in Table VI. It is observed that the model’s accuracy in “All” classes presents a unimodal distribution with respect to the number of old-class experts. On the two dataset, the performance is highest when the number of new and old experts is equal. This indicates that overemphasizing either new-class or old-class experts negatively affects performance due to the disruption of balance. Additionally, we noted that in both datasets, when the number of old-class experts exceeds that of new-class experts, the accuracy in the “Old” classes remains nearly unchanged or increases, while the accuracy in the “New” classes significantly declines, and vice versa. This aligns with the motivation behind our MEA module.
TABLE VI: Performance of the proposed AdaptGCD equipped with different number of old and new experts on Aircraft and Scars datasets.
$|\mathcal{T}^{old}|$ $|\mathcal{T}^{new}|$ Aircraft Scars
All Old New All Old New
2 6 57.0 63.6 53.7 61.7 76.6 54.6
3 5 57.4 63.9 54.2 62.4 79.7 54.0
4 4 57.9 65.2 54.3 62.7 80.6 54.0
5 3 57.4 64.8 53.7 62.0 79.6 53.5
6 2 55.9 65.1 51.3 61.7 81.1 52.3
V-E Discussion
1) Scalability with DINOv2 Backbone: To validate the scalability of the proposal across different pretrained models, we conduct experiments adapting our AdaptGCD to the DINOv2 [81] backbone, as shown in Table VII. Compared to the advanced method, our proposal achieves an overall performance improvement on the “All” classes across three datasets. These observations further illustrate the high adaptability of the proposed adapter-based approach for different pre-trained models.
TABLE VII: Performance on the semantic shift benchmark with the pretrained DINOv2 model.
Methods Backbone CUB-200 Scars Aircraft
All Old New All Old New All Old New
$k$-means [75] DINOv2 67.6 60.6 71.1 29.4 24.5 31.8 18.9 16.9 19.9
GCD [13] DINOv2 71.9 71.2 72.3 65.7 67.8 64.7 55.4 47.9 59.2
SimGCD [55] DINOv2 71.5 78.1 68.3 71.5 81.9 66.6 63.9 69.9 60.9
$\mu$GCD [49] DINOv2 74.0 75.9 73.1 76.1 91.0 68.9 66.3 68.7 65.1
CiPR [82] DINOv2 78.3 73.4 80.8 66.7 77.0 61.8 - - -
AdaptGCD DINOv2 78.7 82.1 77.0 75.8 85.5 71.1 67.7 73.3 64.9
TABLE VIII: Performance of different PEFT methods on CUB-200 and Scars datasets. #Params indicates the number of tunable parameters in backbone of the method.
Methods #Params CUB-200 Scars
All Old New All Old New
LoRA 0.6M 66.6 74.3 62.7 57.6 73.0 50.2
Side Tuning 0.7M 62.1 67.8 59.3 51.6 73.3 41.1
AdaptGCD (single adapter) 0.6M 67.5 75.3 63.5 59.8 77.4 51.2
LoRA 4.6M 63.9 71.7 60.0 54.5 72.1 46.0
Side Tuning 4.4M 62.2 68.1 59.3 53.8 74.9 43.6
AdaptGCD (MEA) 4.8M 68.8 74.5 65.9 62.7 80.6 54.0
2) Comparison with Different Efficient Tuning Methods: To further illustrate that the performance improvement of the proposed method does not solely stem from the structure of the additional parameters, we further reproduce some other common efficient tuning methods, including LoRA [83] and Side Tuning [84, 85] for comparison with AdaptGCD, as shown in Table VIII. To ensure fairness, these competitors are configured with a similar number of tunable parameters as those of the proposed single adapter and multiple adapter methods, respectively. Compared to LoRA, the proposed method incorporates nonlinear layers, introducing greater flexibility and thus achieving higher performance. In contrast to Side Tuning, The proposal involves slight modifications to the original Transformer rather than training an additional light Transformer, which could be more advantageous with limited labels. These observations further validate the superiority of the proposed AdaptGCD.
3) Visualization of the Attention Maps: To better appreciate the advantages of the AdaptGCD, we extracted the attention maps corresponding to the [CLS] token from different heads in the last block of the ViT, highlighting the top 10 most notable patches in red for each head in Fig. 4. Upon observation, it is evident that compared to SimGCD and DINO, AdaptGCD focuses more on the foreground and exhibits more varied attention regions in different heads (in heads 3/7/11/12), showing the superior adaptability of AdaptGCD to the CUB-200 datasets. It is also worth noting that in regions like heads 3/7/11, SimGCD fails to inherit the correct attention on the foreground from the DINO model, while AdaptGCD refines it based on the pretrained model. These findings further demonstrate the motivation to protect the pretrained information.
Figure 4: Attention visualization for 12 heads in the last blocks of backbone on CUB-200. “Pretrained(DINO)” represents the results from pretrained backbone without any additional training while “SimGCD” and “AdaptGCD” are those from models trained via the corresponding methods.
Figure 5: The confusion maps for the old and new classes for the models trained with and without the route assignment losses $\mathcal{L}_{ra}$ on the semantic shift benchmark.
TABLE IX: Experimental results with different number of category $K$ in the network. The estimated category number is obtained using the off-the-shelf method from [13].
Methods Known $K$ CUB($K$=200) Scars($K$=196)
All Old New All Old New
GCD 51.3 56.6 48.7 39.0 57.6 29.9
SimGCD 60.3 65.6 57.7 53.8 71.9 45.0
SPTNet 64.6 70.5 61.6 59.3 61.8 58.1
AdaptGCD 68.8 74.5 65.9 62.7 80.6 54.0
GCD ✗(w/ Est.) 47.1 55.1 44.8 35.0 56.0 24.8
SimGCD ✗(w/ Est.) 61.5 66.4 59.1 49.1 65.1 41.3
SPTNet ✗(w/ Est.) 65.2 71.0 62.3 - - -
AdaptGCD ✗(w/ Est.) 67.2 74.6 63.5 62.6 80.1 54.1
@@ V-E Discussion
[…]
@@
4) Visualization of the Confusion Maps: In addition, to further investigate the role of the route assignment constraint in mitigating the bias towards old classes, we also visualized the confusion matrix of methods trained with and without the $\mathcal{L}_{ra}$ on the three datasets from semantic shift benchmark in Fig. 5. It is observed that with the constraint $\mathcal{L}_{ra}$, the amount of new-class data misclassified as the old classes decreases, thus substantiating the effectiveness of the constraint in alleviating the preference for the old classes.
5) Scalability under Unknown Category Number: Due to the inaccessibility of the total number of categories (GT) in real-world scenarios, we evaluate the proposed AdaptGCD with the estimated number of categories provided by the off-the-shelf method [13] and a larger number of classes in Table IX. The experiments are conducted in the CUB-200 and Scars dataset, where the ground truth of the total class number $K$ are 200 and 196 and the estimated class number $K_{est}$ given by [13] are 231 and 230. It is observed that the proposed AdaptGCD shows overall superiority under different initialized number of classes, demonstrating robustness with estimation errors.
6) Visualization of Assignment Probability: To verify the role of the route assignment constraint, we also visualize the assignment probabilities for samples predicted as old and new classes under the original SimGCD loss, the loss with $\mathcal{L}_{ba}$, and the loss with $\mathcal{L}_{ba}+\mathcal{L}_{cba}$ on Aircraft dataset in Fig. 6. Without any additional loss, the “winner-takes-all” phenomenon occurs and the imbalance is greatly alleviated via $\mathcal{L}_{ba}$. Note that a significant portion of samples predicted as old or new classes are distributed to their corresponding experts with the addition of $\mathcal{L}_{ba}$ and $\mathcal{L}_{cba}$. It validates the effectiveness of $\mathcal{L}_{cba}$ to separate the data from old and new classes on the route level.
Figure 6: Visualization of the route assignment probability for samples predicted as old and new classes in 8 experts.
Figure 7: $t$-SNE visualization of the proposed methods based on (a) single adapter and (b) multiple adapters. The samples in the old classes are marked in blue while those in new classes are marked in red.
7) Visualization of $t$-SNE Results of the Features: To further validate the impact of AdaptGCD on the representations, $t$-SNE [86] is utilized to visualize the features generated by using a single adapter versus multi-expert adapters (with constraints). As shown in Fig. 7, compared to the single adapter, the features generated by the multiple adapters for the new classes (in red) exhibit less interference with the old classes (in blue) and demonstrate greater compactness. These observations confirm the effectiveness of the multi-expert adapters.
VI Conclusion
@@
In this study, we propose a novel AdaptGCD in generalized category discovery, which is the pioneering work to merge adapter tuning into the GCD framework. It integrates learnable adapters with the fixed backbone to adapt to the downstream GCD task as well as preserves the pretrained priority. Additionally, considering the imbalance between the supervision of the old and new classes, we present a multi-expert adapter structure and a route assignment constraint to separate the data from old and new classes into different expert groups, further reducing mutual interference. The efficacy of the proposed strategy is demonstrated in 7 datasets, including 3 generic datasets 3 fine-grained datasets, and a long-tailed dataset. It can also serve as a plug-and-play module to further improve the performance of advanced GCD methods. Looking forward to future work, a promising direction is to explore more effective fine-tuning strategies in GCD tasks, including the RepAdapter [31], or the combination of different strategies like GLoRA [27]. In terms of limitations, the proposed method introduces additional parameters, which consume extra computational cost during training and inference.
VII Acknowledgement
The authors are thankful for the financial support by the National Natural Science Foundation of China (U22B2048, 62476274, 62206293).